NPV-Mutually exclusive projects???Hook Industries is considering the replacement of one of its old drill presses. Three alternative replacement presses are under consideration. The relevant cash flows associated with each are shown in the following? table:

Initial investment (CF0) Press A Press B Press C
$84,500 $59,500 $129,500
Year Cash inflows (CFt)
1 $17,500 $11,900 $49,600
2 $17,500 $14,200 $30,400
3 $17,500 $16,300 $19,800
4 $17,500 $18,400 $20,200
5 $17,500 $19,500 $20,100
6 $17,500 $25,300 $29,800
7 $17,500 - $40,200
8 $17,500 - $49,800

The? firm's cost of capital is 15?%.

a.??Calculate the net present value ?(NPV?) of each press.

b.??Using? NPV, evaluate the acceptability of each press.

c.??Rank the presses from best to worst using NPV.

d.??Calculate the profitability index? (PI) for each press.

e.??Rank the presses from best to worst using PI.

(Round to the nearest cent for all)
Answer #1

Solution a.

NPV (Press A) = -84500 + 17500 * ((1-(1+15%)^(-8))/15%) = -$5971.87

NPV (Press B) = -59500 + 11900/(1+15%) + 14200/(1+15%)^2 + 16300/(1+15%)^3 + 18400/(1+15%)^4 + 19500/(1+15%)^5 + 25300/(1+15%)^6 = $3455.67

NPV (Press C) = -129500 + 49600/(1+15%) + 30400/(1+15%)^2 + 19800/(1+15%)^3 + 20200/(1+15%)^4 + 20100/(1+15%)^5 + 29800/(1+15%)^6 + 40200/(1+15%)^7 + 49800/(1+15%)^8 = $15454.43

Solution b. Press B & Press C are acceptable as NPV is greater than 0

Solution c. Using NPV, ranking of the press from best to worst is

  1. Press C
  2. Press B
  3. Press A

Solution d.

Profitability Index = (NPV + Initial Investment)/Initial Investment

Profitability Index (Press A) = (-5971.87 + 84500)/84500 = 0.93

Profitability Index (Press B) = (3455.67 + 59500)/59500 = 1.06

Profitability Index (Press C) = (15454.43 + 129500)/129500 = 1.12

Solution e. Using Profitability Index, ranking of the press from best to worst is

  1. Press C
  2. Press B
  3. Press A